OpenStudy (anonymous):

Explain, in complete sentences, how you would divide the following expression: .

OpenStudy (anonymous):

the 12th root of x^5 times the sixth root of x

Parth (parthkohli):

\[\rm x^{\large {5 \over 12}}\times x^{\large{1 \over 6}}\]

Parth (parthkohli):

Add the exponents, as you may have learnt already.

Parth (parthkohli):

But where is the division?

OpenStudy (anonymous):

i dont know that how it was writen and the answer would be \[x ^{\left(\begin{matrix}7 \\ 12\end{matrix}\right)}\]